OpenStudy (anonymous):

Two six-sided dice are rolled (one red and one green). Some possibilities are (Red=1,Green=5) or (Red=2,Green=2) etc. (a) How many total possibilities are there? For the rest of the questions, we will assume that the dice are fair and that all of the possibilities in (a) are equally likely. (b) What is the probability that the sum on the two dice comes out to be 9? (c) What is the probability that the sum on the two dice comes out to be 12? (d) What is the probability that the numbers on the two dice are equal?

OpenStudy (mathmale):

We're throwing two dice at the same time. Does the throw of one die affect the throw of the other die? Throwing either die first could produce any of 6 possible outcomes, all of which are equally likely because the dice are fair. Throwing the other die also produces any of 6 possible outcomes. Assuming that tossing each die is independent of tossing the other one, how many possible outcomes could there be?

OpenStudy (anonymous):

6*6 = 36

OpenStudy (anonymous):

Right? @mathmale

OpenStudy (anonymous):

I got c also it is 1/36

OpenStudy (anonymous):

How about b and d? @mathmale

OpenStudy (mathmale):

Let's state the "answers" accurately. The number of possible outcomes, when two dice are thrown at the same time, is .... ?

OpenStudy (anonymous):

36?

OpenStudy (mathmale):

Yes. If you answer in complete sentences, your answers are likely to have more meaning for you as well as for anyone who reads your work. Now, for part (b):

OpenStudy (mathmale):

Write out the actual outcomes of tossing two dice together that would result in a sum of 9.

OpenStudy (anonymous):

Okay For b part, is the probability that the sum on the two dice comes out to be 9 is 4/36 or 1/9?

OpenStudy (mathmale):

Mind writing out all of the possibilities of what the 2 dice would look like? For example: R:4 with G: 5 (since 4+5 = 9).

OpenStudy (anonymous):

Yes 3+6, 4+5, 5+4 and 6+3

OpenStudy (anonymous):

Right so 4/36 or 1/9?

OpenStudy (anonymous):

@mathmale please help:(

OpenStudy (anonymous):

@nkaty You are right. 4/36 = 1/9 is correct for (b)

OpenStudy (anonymous):

And for (d), there are obviously 6 cases wherein the dice fall with the same face up, right? So you should be able to calculate the probability for the event in (d).